# SITransformer: Shared Information-Guided Transformer for Extreme Multimodal Summarization

### v1

@@ Abstract.
Extreme Multimodal Summarization with Multimodal Output (XMSMO) becomes an attractive summarization approach by integrating various types of information to create extremely concise yet informative summaries for individual modalities. Existing methods overlook the issue that multimodal data often contains more topic irrelevant information, which can mislead the model into producing inaccurate summaries especially for extremely short ones. In this paper, we propose SITransformer, a Shared Information-guided Transformer for extreme multimodal summarization. It has a shared information guided pipeline which involves a cross-modal shared information extractor and a cross-modal interaction module. The extractor formulates semantically shared salient information from different modalities by devising a novel filtering process consisting of a differentiable top-k selector and a shared-information guided gating unit. As a result, the common, salient, and relevant contents across modalities are identified. Next, a transformer with cross-modal attentions is developed for intra- and inter-modality learning with the shared information guidance to produce the extreme summary. Comprehensive experiments demonstrate that SITransformer significantly enhances the summarization quality for both video and text summaries for XMSMO. Our code will be publicly available at https://github.com/SichengLeoLiu/MMAsia24-XMSMO.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Video summarizationccs: Computing methodologies Information extraction
@@ 1. Introduction
In the digital age, information overload is a significant challenge due to the rapid production of data that surpasses human processing abilities. Summarization techniques aim to distill essential information quickly, aiding in managing this overload. Traditional text summarization and video summarization methods have been extensively studied (Kupiec et al., 1995; Liu and Lapata, 2019; Zhang et al., 2016). However, with the rise of multimodal content on the Internet, conventional unimodal summarization methods are increasingly inadequate. Hence, multimodal summarization with multimodal output (MSMO) (Zhu et al., 2018), which integrates information from multiple modalities, has emerged to address this gap. By incorporating diverse data sources such as videos and text documents, with advanced deep learning techniques (Vaswani et al., 2017; Hu et al., 2019), it enhances summarization quality and user experience. To further simplify and expedite information acquisition, extreme summarization was introduced into the MSMO task (Tang et al., 2023b, a), proposing eXtreme Multimodal Summarization with Multimodal Output (XMSMO). XMSMO condenses a video-document pair into a single frame and a single sentence. This extremely concise multimodal output allows users to quickly grasp the essence and decide whether to engage further.
@@
To address this issue, we posit that shared information across different modalities is capable to provide critical insights for comprehending the topic, thereby facilitating better utilization of the relevant information from both shared and non-shared contents for the summary. Thus, this paper proposes an innovative multimodal summarization method called SITransformer for the video-document XMSMO task. Overall, SITransformer consists of a cross-modal share information extractor and a cross-modal interaction module for the effective exploitation and utilization of shared information guidance. Specifically, the extractor aims to extract the salient information shared across modalities by a differentiable top-k selection, which ranks and selects the most salient features across modalities. The shared information is then used to control the information gate units, which determine which portions of the source features should be retained and which should be discarded, thereby filtering out features irrelevant to the topic and summarization. The cross-modal interaction module then processes the extracted relevant information using visual and textual transformers with cross-modality attentions. This enables both unimodal and cross-modal information learning to produce the final summarization.
@@
    Comprehensive experiments on a large-scale multimodal dataset demonstrates the effectiveness of SITransformer in achieving state-of-the-art performance.
@@ 2. Related Work
Unimodal Summarization focuses on summarizing content from a single modality, such as text or video. In text summarization, extractive methods select key sentences from the original document (Liu, 2019). Abstractive methods instead generate new sentences, often using Transformer-based models like BART (Lewis et al., 2019). Hybrid methods, such as the Pointer-Generator Network (See et al., 2017; Tang et al., 2022), combine extraction and generation of content for the final summary. Video summarization involves selecting informative frames to represent the main content. Early studies relied on methods such as LSTMs to formulate temporal patterns (Zhang et al., 2016), while recent methods used Transformer-based methods to capture frame relationships (Zhao et al., 2022). These unimodal methods establish a critical foundation for multimodal summarization.
@@
Extreme summarization condenses original data into extremely concise content. For instance, extreme video summarization aims to distill a video into a single frame, while extreme text summarization focuses on summarizing a document into one sentence. Tang et al. (2023b, a) introduced the XMSMO task, which summarizes a video-document pair into a single sentence and a single frame. Their hierarchical model integrated information across modalities at sentence-scene and video-document levels. By minimizing Wasserstein distances between selected frames/texts and original videos/documents respectively, the model improves summarization coherence. These studies offer a feasible approach for extreme summarization but are often affected by irrelevant information in rich multimodal data, resulting in less decent summarization quality.
@@ 3. Methodology
The overall pipeline of the proposed SITransformer is depicted in Figure 2. It consists of four core stages with the shared salient information guidance. Firstly, a multimodal embedding module, which consists of multimodal encoders, is used to embed the input video and document into latent features. Next, a cross-modal shared information extractor processes the latent features with a noise filtering process. A cross-modal interaction module further guides the summarization of each modality with the extracted shared information. Finally, an extreme summarization decoder creates the corresponding extreme summaries for video and document.
@@ 3.2. Multimodal Embedding Module
To achieve a basic alignment between visual and textual modalities, we apply the widely used multimodal embedding model, CLIP (Radford et al., 2021), as our latent feature encoder. The CLIP’s multimodal latent space enables the visual and textual contents with related semantics to share similar embeddings, which is able to facilitate the cross-modal shared information exploration.
@@ 3.3. Cross-Modal Shared Information Extractor
The low and high-level features contain rich information relevant to the major topic. However, they often include a significant amount of irrelevant information, making precise summarization challenging. To reduce such negative impact, a cross-modal shared information extractor with noise filtering mechanism is devised. The core idea is to leverage shared salient information, presenting in both modalities, to regulate information gates and filter out irrelevant information from the shared and non-shared feature embeddings accordingly.
@@
The noise filtering mechanism is designed with a Differentiable Top-k (NFDT) (Cordonnier et al., 2021) strategy as depicted in Figure 3. Specifically, linear layers are first adopted to obtain salient information scores for each word and frame:
@@
Next, Gumbel softmax functions (Jang et al., 2016) are utilized to standardize the score between 0 and 1. Compared to the regular softmax function, the Gumbel softmax function is differentiable by integrating noise sampled from Gumbel distribution. Mathematically, we first sample noise $g$ by:
@@
We can obtain the score list $\mathbf{S}_{T^{\prime}}=\{score_{t,i}|i\in[1,n]\}$ and $\mathbf{S}_{V^{\prime}}=\{score_{v,j}|j\in[1,m]\}$ for all individual texts and frames. The top-k features with highest estimated scores are selected to construct the potential multimodal shared salient information $\mathbf{X}_{S}$ as:
@@ 3.6. Extreme Summarization Decoder
The cover frame selection can be regarded as a classification task. We use a linear layer with softmax activation to score the frames and identify the one $\mathbf{O}_{V}$ with the maximum score as the cover frame:
@@ 3.7. Loss and Optimization
$\mathcal{L}_{T}$ calculates the distributional distance between the one-sentence-summary and the input document, where $p_{t}$ and $q_{t}$ are probability mass functions of document feature $\mathbf{X}^{H}_{T}$ and summary feature $\mathbf{X}_{TO}$. Similarly, $\mathcal{L}_{V}$ represents the distance between cover-frame summary and the input video probability mass function. $\mathcal{L}_{O}$ further ensures the output summary is consistent between sentences and frames.
@@
where $P_{LM}(\mathbf{O}_{T})$ is the probability of the one-sentence assigned by a large language model (GPT-2 in practice). $P_{U}(\mathbf{O}_{T})$ is the unigram probability, which can be calculated as:
(20) $P_{U}(\mathbf{O}_{T})=\prod_{t\in\mathbf{O}_{T}}P(t).$
To this end, the overall loss can be calculated as:
@@
where $\lambda_{T}$, $\lambda_{V}$, $\lambda_{S}$ and $\lambda_{f}$ are hyper-parameters that used to control the importance of each loss term.
@@ 4.1. Experimental Settings
Evaluation Metrics. Following existing practice (Tang et al., 2023b), for the quantitative evaluation of video summarization, intersection over union (IoU) (Sharghi et al., 2017) ad frame accuracy (FA) (Messaoud et al., 2021) are adapted. In addition, ROUGE (Lin, 2004) metric is adopted for the text summarization.
@@ 4.2. Quantitative Analysis
Table 1. Quantitative evaluation on the multimodal BBC news dataset (Tang et al., 2023b).
@@ 4.4. Ablation Study
We conducted ablation studies to verify the effectiveness of each proposed components. The experimental results are listed in Table 1. Firstly, we removed the shared information selection process in obtaining the multimodal shared salient information. The significantly degraded performance indicates the importance of extracting shared salient information to guide the summarization. We also assess the impacts of Gumbel-softmax by replacing with standard softmax, and evaluate shared information guided gate design by its removal in our proposed NFDT. The results demonstrate their positive contributions to enhancing the extraction of shared salient information. Lastly, to verify the effectiveness of the proposed noise filter design, we replace our NFDT noise filter with two different types of filters, namely random feature selector and noise filter with cosine similarity (NFCS). The former filters shared salient features by random selection while the later applies cosine similarity based ranking for salient information identification. Compared to our best model, both visual and textual summary metrics degrade, demonstrating the performance of our proposed novel noise filter.
Figure 5. Impacts of shared salient information amount in percentage on quantitative metrics
@@ 4.5. Impacts of Shared Information Amount
The amount of shared salient information to extract is an important hyper-parameters. We visualize its impacts during the multimodal shared information extraction process in Figure  5. It shows that retaining 50% of shared information leads to better performance in terms of most metrics except IoU. Selecting either an excessive or insufficient number of features can adversely affect performance. Choosing too many shared salient features may result in the inclusion of undesirable, irrelevant information, leading to the loss of valuable data. Conversely, using too few shared salient features may fail to adequately control the noise filter, also causing performance degradation.

### v3

@@ Abstract.
Extreme Multimodal Summarization with Multimodal Output (XMSMO) becomes an attractive summarization approach by integrating various types of information to create extremely concise yet informative summaries for individual modalities. Existing methods overlook the issue that multimodal data often contains more topic irrelevant information, which can mislead the model into producing inaccurate summaries especially for extremely short ones. In this paper, we propose SITransformer, a Shared Information-guided Transformer for extreme multimodal summarization. It has a shared information guided pipeline which involves a cross-modal shared information extractor and a cross-modal interaction module. The extractor formulates semantically shared salient information from different modalities by devising a novel filtering process consisting of a differentiable top-k selector and a shared-information guided gating unit. As a result, the common, salient, and relevant contents across modalities are identified. Next, a transformer with cross-modal attentions is developed for intra- and inter-modality learning with the shared information guidance to produce the extreme summary. Comprehensive experiments demonstrate that SITransformer significantly enhances the summarization quality for both video and text summaries for XMSMO. Our code is publicly available at https://github.com/SichengLeoLiu/MMAsia24-XMSMO.
@@
journalyear: 2024copyright: rightsretainedconference: ACM Multimedia Asia; December 3–6, 2024; Auckland, New Zealandbooktitle: ACM Multimedia Asia (MMASIA ’24), December 3–6, 2024, Auckland, New Zealanddoi: 10.1145/3696409.3700234isbn: 979-8-4007-1273-9/24/12ccs: Computing methodologies Video summarizationccs: Computing methodologies Information extraction
@@ 1. Introduction
In the digital age, information overload is a major issue due to the rapid production of data that surpasses human processing abilities. Summarization techniques aim to distill essential information quickly, aiding in managing this overload. Traditional text summarization and video summarization methods have been extensively studied (Kupiec et al., 1995; Liu and Lapata, 2019; Zhang et al., 2016). However, with the rise of multimodal content on the Internet, conventional unimodal summarization methods are increasingly inadequate. Hence, multimodal summarization with multimodal output (MSMO) (Zhu et al., 2018), which integrates information from multiple modalities, has emerged to address this gap. By incorporating diverse data sources such as videos and text documents, with advanced deep learning techniques (Vaswani et al., 2017; Hu et al., 2019), it enhances summarization quality and user experience. To further simplify and expedite information acquisition, extreme summarization was introduced into the MSMO task (Tang et al., 2023b, a), proposing eXtreme Multimodal Summarization with Multimodal Output (XMSMO). XMSMO condenses a video-document pair into a single frame and a single sentence. This extremely concise multimodal output allows users to quickly grasp the essence and decide whether to engage further.
@@
To address this issue, we posit that shared information across different modalities is able to provide critical insights for comprehending the topic, thereby facilitating better utilization of the relevant information from both shared and non-shared contents for the summary. Thus, this paper proposes an innovative multimodal summarization method called SITransformer for the video-document XMSMO task. Overall, SITransformer consists of a cross-modal share information extractor and a cross-modal interaction module for the effective exploitation and utilization of shared information guidance. Specifically, the extractor aims to extract the salient information shared across modalities by a differentiable top-k selection, which ranks and selects the most salient features across modalities. The shared information is then used to control the information gate units, which determine which portions of the source features should be retained and which should be discarded, thereby filtering out features irrelevant to the topic and summarization. The cross-modal interaction module then processes the extracted relevant information using visual and textual transformers with cross-modality attentions. This enables both unimodal and cross-modal information learning to produce the final summarization.
@@
    Comprehensive experiments on a large-scale multimodal dataset demonstrate the effectiveness of SITransformer in achieving state-of-the-art performance.
@@ 2. Related Work
Unimodal Summarization focuses on summarizing content from a single modality, such as text or video. In text summarization, extractive methods select key sentences from the original document (Liu, 2019). Abstractive methods instead generate new sentences, often using Transformer-based models like BART (Lewis et al., 2019). Hybrid methods, such as the Pointer-Generator Network (See et al., 2017; Tang et al., 2022), combine extraction and generation of content for the final summary. More recently, large language model-based methods for text summarization have been explored, but they are more prone to hallucination and face challenges in fine-tuning for domain-specific knowledge compared to traditional methods (Jin et al., 2024). Video summarization involves selecting informative frames to represent the main content. Early studies relied on methods such as LSTMs to formulate temporal patterns (Zhang et al., 2016), while recent methods used Transformer-based methods to capture frame relationships (Zhao et al., 2022). These unimodal methods establish a critical foundation for multimodal summarization.
@@
Extreme summarization condenses original data into extremely concise content. For instance, extreme video summarization aims to distill a video into a single frame, while extreme text summarization focuses on summarizing a document into one sentence. Tang et al. (2023b, a) introduced the XMSMO task, which summarizes a video-document pair into a single sentence and a single frame. Their hierarchical model integrated information across modalities at sentence-scene and video-document levels. By minimizing Wasserstein distances between selected frames/texts and original videos/documents respectively, the model improves summarization coherence. These studies offer a feasible approach for extreme summarization but are often affected by irrelevant information in rich multimodal data, resulting in less promising summarization quality.
@@ 3. Methodology
The proposed SITransformer pipeline, shown in Figure 2, includes four core stages with shared salient information guidance.
@@ 3.2. Multimodal Embedding Module
To achieve an initial alignment between visual and textual modalities, we apply the widely used multimodal embedding model, CLIP (Radford et al., 2021), as our latent feature encoder. The CLIP’s multimodal latent space enables the visual and textual contents with related semantics to share similar embeddings, which is able to facilitate the cross-modal shared information exploration.
@@ 3.3. Cross-Modal Shared Information Extractor
The low and high-level features contain rich information relevant to the major topic. However, they often include significant irrelevant information, making precise summarization challenging. To reduce such negative impact, a cross-modal shared information extractor with noise filtering mechanism is devised. The core idea is to leverage shared salient information, presenting in both modalities, to regulate information gates and filter out irrelevant information from the shared and non-shared feature embeddings accordingly.
@@
The noise filtering mechanism is designed with a Differentiable Top-k (NFDT) (Cordonnier et al., 2021) strategy as depicted in Figure 3. Specifically, linear layers are first adopted to obtain word/frame salient scores:
@@
Next, Gumbel softmax functions (Jang et al., 2016) are utilized to standardize the score between 0 and 1. Compared to the regular softmax function, the Gumbel softmax function is differentiable by integrating noise $g$ sampled from Gumbel distribution by:
@@
We can obtain the score list $\mathbf{S}_{T^{\prime}}=\{score_{t,i}|i\in[1,n]\}$ and $\mathbf{S}_{V^{\prime}}=\{score_{v,j}|j\in[1,m]\}$ for all individual texts and frames. The top-k features with the highest estimated scores are selected to construct the potential multimodal shared salient information $\mathbf{X}_{S}$:
@@ 3.6. Extreme Summarization Decoder
The cover frame selection is treated as a classification task. We use a linear layer with softmax activation to score the frames and identify the one $\mathbf{O}_{V}$ with the maximum score as the cover frame:
@@ 3.7. Loss and Optimization
$\mathcal{L}_{T}$ calculates the distributional distance between the one-sentence-summary and the input document, where $p_{t}$ and $q_{t}$ are probability mass functions of document feature $\mathbf{X}^{H}_{T}$ and summary feature $\mathbf{X}_{TO}$. Similarly, $\mathcal{L}_{V}$ represents the distance between the cover-frame summary and the input video probability mass function. $\mathcal{L}_{O}$ ensures the output summary is consistent between sentences and frames.
@@
where $P_{LM}(\mathbf{O}_{T})$ is the probability of the one-sentence assigned by a large language model (GPT-2 in practice) and $P_{U}(\mathbf{O}_{T})$ is the unigram probability. To this end, the overall loss can be calculated as:
@@
where $\lambda_{T}$, $\lambda_{V}$, $\lambda_{S}$, $\lambda_{f}$ are hyper-parameters to control loss weight.
@@ 4.1. Experimental Settings
Evaluation Metrics. Following existing practice (Tang et al., 2023b), for the quantitative evaluation of video summarization, intersection over union (IoU) (Sharghi et al., 2017) and frame accuracy (FA) (Messaoud et al., 2021) are adapted. In addition, ROUGE (Lin, 2004) metric is adopted for the text summarization.
@@ 4.2. Quantitative Analysis
Table 1. Quantitative evaluation on BBC news dataset (Tang et al., 2023b).
@@ 4.4. Ablation Study
We conducted ablation studies to verify the effectiveness of each proposed component. The experimental results are listed in Table 1. Firstly, we removed the shared information selection process in obtaining the multimodal shared salient information. The significantly degraded performance indicates the importance of extracting shared salient information to guide the summarization. We also assess the impacts of Gumbel-softmax by replacing with standard softmax, and evaluate shared information guided gate design by its removal in our proposed NFDT. The results demonstrate their positive contributions to enhancing the extraction of shared salient information. Lastly, to verify the effectiveness of the proposed noise filter design, we replace our NFDT noise filter with two different types of filters, namely random feature selector and noise filter with cosine similarity (NFCS). The former filters shared salient features by random selection while the later applies cosine similarity based ranking for salient information identification. Compared to our best model, both visual and textual summary metrics degrade, demonstrating the performance of our proposed novel noise filter.
Figure 5. Impacts of shared salient information amount
@@ 4.5. Impacts of Shared Information Amount
The amount of shared salient information to extract is an important hyper-parameters. We visualize its impacts during the multimodal shared information extraction process in Figure  5. It shows that retaining 50% of shared information improves most metrics except IoU. Selecting either an excessive or insufficient number of features can adversely affect performance. Choosing too many shared salient features may result in the inclusion of undesirable, irrelevant information, leading to the loss of valuable data. Conversely, using too few shared salient features may fail to adequately control the noise filter, also causing performance degradation.